OpenStudy (anonymous):

In a 45°-45°- 90° triangle, if the legs each have a length of 4, then what is the length of the hypotenuse? A. the square root of2 B. 4the square root of2 C. 4 D. 4the square root of3

OpenStudy (accessdenied):

Do you know the ratio of the side-lengths to the hypotenuse? Otherwise we could also use Pythagorean theorem if we had forgotten.

OpenStudy (anonymous):

I just needed this one explained.. I did not understand how to do special right triangles.

OpenStudy (accessdenied):

On one hand, we have some specific lengths to memorize. For any 45-45-90 triangle, there is a specific proportion relating the length of the legs to the length of the hypotenuse. It is normally a task of simply memorizing those ratios to make this very painless. But it all rolls back to the Pythagorean theorem for proof methods. Say we had a leg side length of s and we call the hypotenuse h. Then Pythagorean thm tells us that: s^2 + s^2 = h^2.  We simplify by adding the like terms: 2s^2 = h^2 And then the square root of both sides gives us: sqrt(2) s = h. This general statement always holds true given a leg side-length s or a hypotenuse length h, for all 45-45-90 triangles... makes sense so far?

OpenStudy (accessdenied):

The reason we really have the special triangles, though, is for a simplification reason. I showed that sqrt(2) * s = h for all 45-45-90 triangles. If we simply know this to be true, we erase some work in the process. Instead of writing out the Pythagorean thm, we just say "We know we have a 45-45-90 triangle. And the side length is 4. Now we just plug it in: h = 4 * sqrt(2) "

OpenStudy (accessdenied):

Just like if we had the hypotenuse already, say h = 45 sqrt(2), the equation also holds: 45 sqrt(2) = s * sqrt(2) 45 = s

OpenStudy (anonymous):

So in other words, without knowing the hypotenuse, it would be 4 sqrt(2)?

OpenStudy (accessdenied):

Yes. In general, given a leg side-length of the 45-45-90 triangle, we simply multiply by sqrt(2) to obtain our hypotenuse. And given the hypotenuse, the reverse direction of dividing by sqrt(2) gets us back to side length.

OpenStudy (anonymous):

So final answer would be sqrt(2)?

OpenStudy (accessdenied):

We already had the hypotenuse as 4sqrt(2). What did you do to get sqrt(2) only?

OpenStudy (anonymous):

Idk.. sorry lol

OpenStudy (accessdenied):

No problem. :p

OpenStudy (anonymous):

Could you show me how to find X on a triangle from a special right 45, and 30? Ive got a few pics

OpenStudy (accessdenied):

For the first one. 9 is in the same place as which part (s, s sqrt(3), or 2s) in the 30-60-90 triangle I drew? And then for x?

OpenStudy (anonymous):

The s.

OpenStudy (accessdenied):

Hang on...

OpenStudy (anonymous):

And the x is the hypotenuse, SQRT(2)

OpenStudy (anonymous):

you scribbling that out confuses me..lol

OpenStudy (accessdenied):

Okay, let me redraw that.